OpenStudy (anonymous):

find two supplementary angles such that the measure of the first angle is 46 degrees less than three times the measure of the second. first and and second angle is?

OpenStudy (radicalsniper):

X + Y = 180 degrees , X = first angle, Y = second angle. X= 3Y - 46. 3Y - 46 + Y = 180

OpenStudy (anonymous):

4y-46=180?

OpenStudy (anonymous):

4y=226

OpenStudy (anonymous):

y=56.5?

OpenStudy (anonymous):

so first angle is102.5 and second is 56.5?

OpenStudy (radicalsniper):

When you get Y, substitute it in: X = 3Y - 46

OpenStudy (anonymous):

123.5 but where does x fit into?

OpenStudy (radicalsniper):

X is the first angle.

OpenStudy (anonymous):

and to get 2nd its 123.5-46?

OpenStudy (radicalsniper):

You already have the second angle which is Y.

OpenStudy (anonymous):

i see thanks i probably need to study this more then

OpenStudy (radicalsniper):

To check, both angles must be equal to 180 :)